# An automated occultation network for gravitational mapping of the trans-neptunian solar system

### v1

@@ 3.5 Results: Trojans
We therefore investigate changing the telescope pointing from $D=2$ km for the MBAs to 5 km for the Trojans, which will increase the number of observable occultations for Trojans with $d_{T}>2$ km—which is, in fact, all of the LSST-discoverable sources. Figures 4 illustrate the result of a configuration near the new optimum for a \$15M nominal cost. The number and size of telescopes resemble the MBA optimum, but spread out over 1000 km. This wider-spaced array might be harder to maintain; it would, however, also be able to conduct the MBA survey, with the caveat that many forecasted MBA occultations would be “duds” that pass between stations, leading to a higher workload for the array.
The scaling relations of the optimal $d_{T}$ and of $\sigma_{\rm tot}$ with the cost, duration, and per-station costs remain essentially the same as the MBA case. Most of the occultations still occur in low-background situations. Some practical differences are, however, that
@@ 4.3 Dynamical model
There are two relevant non-gravitational forces that act significantly on test particles (the observed MBAs and Trojans)—solar radiation pressure (SRP) and the Yarkovski effect. The accelerations attributable to these are parallel and perpendicular, respectively, to the sun-asteroid vector, lying in the orbital plane, [^6] There is a less important component perpendicular to the orbital plane. and each is inversely proportional to the body diameter, such that
@@
In our previous paper (Gomes et al., 2023), we modeled each of the $k$ parameters as global scaling factors. We improve this model by allowing each asteroid to have a free Yarkovski tangential acceleration parameter. This change is required because our predicted per-object center-of-mass uncertainties are smaller than expected displacements from the Yarkovski force, the latter being of the order of a few kilometers for most MBA targets during a 10 year arc. Therefore, the variance in Yarkovski displacements due to different asteroid shapes and albedos will exceed the other measurement errors. Adding individual Yarkovski accelerations for each asteroid as constrainable parameters solves this issue, with the only downside that each orbit solution needs an extra data point before it starts being useful for tidal field information: four initial occultations are needed instead of three. On the positive side, our occultation survey would successfully detect this Yarkovski acceleration parameter on most of our targets, an achievement that otherwise would require many decades of observations (Hung et al., 2023). For SRP, which has a smaller contribution to orbital displacement, we keep a single global parameter $k_{\text{SRP}}.$
The final element of the dynamical model is a set of 5 point masses at $400$ AU and different fixed positions. Gomes et al. (2023) show that a Planet X tidal field at any location in the sky can be expressed as a linear combination of the tidal fields from these 5 putative masses. Hence forecasting the covariance matrix of these 5 mass parameters is sufficient to place bounds on the mass of a Planet X in any chosen sky direction. The quadrupole moment of the Kuiper belt mass distribution generates a gravitational field on the Trojans and MBAs that is indistinguishable from a Planet X tidal field. We therefore add to the $5\times 5$ covariance matrix an additional covariance arising from the positions and mass uncertainties of the largest individual TNOs, as well as quadrupole fields from the remaining TNO population, as modeled by the CFEPS project (Kavelaars et al., 2009; Petit et al., 2011; Gladman et al., 2012). This calculation is described in more detail on our previous work (Gomes et al., 2023). The resultant “KBO floor” is dominated by the mass uncertainties of the largest moonless TNOs.
@@
The remaining model elements—ring potentials, solar $J_{2}$, the mass of Pluto, and tangential acceleration due to Yarkovski effect—are added one by one to the simulations with two slightly different parameter values, so that their derivatives can be obtained through finite differences. To implement the ring potentials and non-gravitational forces we make use of the REBOUNDx package (Tamayo et al., 2019), as described in Gomes et al. (2023). Derivatives with respect to the solar radiation pressure parameter are obtained by rescaling the derivatives with respect to the solar mass.
@@ 4.4 Fisher uncertainties
Once equipped with our list of simulated observations and their uncertainties, as well as derivatives of model parameters with respect to the observables, we compute the Fisher Matrix to estimate lower bounds on the uncertainties $\sigma_{M}$ in the mass of a hypothetical Planet X at 400 AU. We follow the method delineated in Gomes et al. (2023), where a Fisher Matrix with all global parameters is first calculated for the set of existing historical observations of the major planets. The observations we use are the same ones enumerated in the aforementioned paper and they consist of astrometry and radio ranging measurements. Then, for each test particle, the matrix is augmented with its initial-state-vector parameters and simulated observations, and subsequently marginalized over the state vector. The Fisher matrix thus remains fixed in size even as we accumulate constraints from millions of test bodies. Finally, we invert the total Fisher matrix to form a covariance matrix, and extract the $5\times 5$ sub-matrix corresponding to the five putative distant point masses. To the resultant covariance matrix, we add the previously mentioned Kuiper Belt floor covariance. Then, for any given position in the sky, a transformation of the covariance matrix, as detailed in Appendix A of Gomes et al. (2023), yields us the lower bound uncertainty $\sigma_{M}$ on the mass of a Planet X in any chosen direction from the Sun.
@@ 5 Simulated constraints
Three scenarios are presented in Figure 6. The first one (top left), assumes the array would be exclusively devoted to MBA occultations. In this case, distance between telescopes is set at $D=2$ km. There is no change in $\sigma_{M}$ at larger $D$, since the majority of occulters are smaller than 2 km and the number of observed occultations becomes independent of $D$. The second scenario (top right) assumes an exclusive Jovian Trojan survey, and sets the distance between telescopes to $D=5$ km. Larger $D$ is favored for the Trojan survey, since they are twice as far as MBAs, making the smallest ones detectable by LSST larger than for MBAs. The third scenario (bottom) assumes a joint MBA/Trojan survey, with $D=5$ km again since the Trojans benefit and the MBA observations are not harmed by increased $D$.
@@ 6 Conclusions
The constraint on missing mass in the solar system would not be the only science benefit of the small-telescope linear array. Our survey would allow the study of other perturbations to asteroid orbits, constraining Yarkovski acceleration parameters and detecting mutual deflections by close asteroid encounters. These data, therefore, would contribute to a better understanding of asteroid densities, albedos, and thermal properties. Such an array would also provide detailed shapes for $\sim 1800$ TNOs, characterizing the distribution of sizes, shapes, albedos, and contact-binary rates of TNOs in many different dynamical states. No other envisioned observation could yield such information for a large sample of pristine products of planetesimal formation.
@@ Appendix B Measurement errors on chord midpoints
where we have defined $B=n_{b}/n_{\star}$ as the background ratio. In the second line, we have exploited the fact that the Fisher information in a Poisson distribution with a mean of $\lambda(q)$ for some parameter $q$ is equal to $(d\lambda/dq)^{2}/\lambda,$ and we can divide the light curve into statistically independent time segments of duration $\Delta t$ with expected mean count rates of $\lambda=n(t)\,\Delta t.$

### v2

@@ 3.5 Results: Trojans
We therefore investigate changing the telescope spacing from $D=2$ km for the MBAs to 5 km for the Trojans, which will increase the number of observable occultations for Trojans with $d_{T}>2$ km—which is, in fact, all of the LSST-discoverable sources. Figures 4 illustrate the result of a configuration near the new optimum for a \$15M nominal cost. The number and size of telescopes resemble the MBA optimum, but spread out over 1000 km. This wider-spaced array might be harder to maintain; it would, however, also be able to conduct the MBA survey, with the caveat that many forecasted MBA occultations would be “duds” that pass between stations, leading to a higher workload for the array.
The scaling relations of the optimal $d_{T}$ and of $\sigma_{\rm tot}$ vs. the cost, duration, and per-station costs remain essentially the same as the MBA case. Most of the occultations still occur in low-background situations. Some practical differences are, however, that
@@ 4.2.1 Shape noise
Additional noise will be generated if a binary system is mistakenly assumed to be a single object, because only one member of the system is detected in occultation. In this case, our best center of mass prediction will be offset by a factor dependent on the binary mass ratio and separation. Near-Earth orbit observations indicate a binary fraction of  15% (Pravec et al., 2006). Assuming a similar figure for the MBAs, we look at the distribution of mass ratios and separations of known km-sized MBA binaries and we find that most have center of mass offsets smaller or comparable to our nominal shape noise error. We, therefore, do not model this additional error, and introduce the following caveat to our results: ideally, the survey should extend until enough detections of each target are made that additional scatter caused by a binary companion could be statistically inferred.
@@ 4.3 Dynamical model
There are two relevant non-gravitational forces that act significantly on test particles (the observed MBAs and Trojans)—solar radiation pressure (SRP) and the Yarkovsky effect. The accelerations attributable to these are parallel and perpendicular, respectively, to the sun-asteroid vector, lying in the orbital plane, [^6] There is a less important component perpendicular to the orbital plane. and each is inversely proportional to the body diameter, such that
@@
In our previous paper (Gomes et al., 2023), we modeled each of the $k$ parameters as global scaling factors. We improve this model by allowing each asteroid to have a free Yarkovsky tangential acceleration parameter. This change is required because our predicted per-object center-of-mass uncertainties are smaller than expected displacements from the Yarkovsky force, the latter being of the order of a few kilometers for most MBA targets during a 10 year arc. Therefore, the variance in Yarkovsky displacements due to different asteroid shapes and albedos will exceed the other measurement errors. Adding individual Yarkovsky accelerations for each asteroid as constrainable parameters solves this issue, with the only downside that each orbit solution needs an extra data point before it starts being useful for tidal field information: four initial occultations are needed instead of three. On the positive side, our occultation survey would successfully detect this Yarkovsky acceleration parameter on most of our targets, an achievement that otherwise would require many decades of observations (Hung et al., 2023). For SRP, which has a smaller contribution to orbital displacement, we keep a single global parameter $k_{\text{SRP}}.$
The final element of the dynamical model is a set of 5 point masses at $400$ AU and different fixed positions. Gomes et al. (2023) show that a Planet X tidal field at any location in the sky can be expressed as a linear combination of the tidal fields from these 5 putative masses. Hence forecasting the covariance matrix of these 5 mass parameters is sufficient to place bounds on the mass of a Planet X in any chosen sky direction. The quadrupole moment of the Kuiper belt mass distribution generates a gravitational field on the Trojans and MBAs that is indistinguishable from a Planet X tidal field. We therefore add to the $5\times 5$ covariance matrix an additional covariance arising from the positions and mass uncertainties of the largest individual TNOs, as well as quadrupole fields from the remaining TNO population, as modeled by the CFEPS project (Kavelaars et al., 2009; Petit et al., 2011; Gladman et al., 2012). This calculation is described in more detail on our previous work (Gomes et al., 2023). The resultant “Kuiper Belt object (KBO) floor” is dominated by the mass uncertainties of the largest moonless TNOs.
@@
The remaining model elements—ring potentials, solar $J_{2}$, the mass of Pluto, and tangential acceleration due to Yarkovsky effect—are added one by one to the simulations with two slightly different parameter values, so that their derivatives can be obtained through finite differences. To implement the ring potentials and non-gravitational forces we make use of the REBOUNDx package (Tamayo et al., 2019), as described in Gomes et al. (2023). Derivatives with respect to the solar radiation pressure parameter are obtained by rescaling the derivatives with respect to the solar mass.
@@ 4.4 Fisher uncertainties
Once equipped with our list of simulated observations and their uncertainties, as well as derivatives of model parameters with respect to the observables, we compute the Fisher Matrix to estimate lower bounds on the uncertainties $\sigma_{M}$ in the mass of a hypothetical Planet X at 400 AU. We follow the method delineated in Gomes et al. (2023): first, a Fisher Matrix is generated for the set of existing observations of the major planets with respect to our global model parameters, namely the masses and state vectors of the planets, masses of $\approx 300$ asteroids, collective mass rings for smaller MBAs and TNOs, and the tidal forces of interest. The observations we use for this step are the same ones enumerated in the aforementioned paper, and they consist of astrometry and radio ranging measurements. Next, for each small-body test particle, the Fisher matrix is augmented with rows and columns for its initial-state-vector parameters (and, in our current case, a Yarkovsky effect strength), and then the Fisher information from its simulated observational data is added in. We can then marginalize the Fisher matrix over this body’s parameters. The Fisher matrix thus remains fixed in size even as we accumulate constraints from millions of test bodies. Finally, we invert the total Fisher matrix to form a covariance matrix, and extract the $5\times 5$ sub-matrix corresponding to the five putative distant point masses. To the resultant covariance matrix, we add the previously mentioned Kuiper Belt floor covariance. Then, for any given position in the sky, a transformation of the covariance matrix, as detailed in Appendix A of Gomes et al. (2023), yields us the lower bound uncertainty $\sigma_{M}$ on the mass of a Planet X in any chosen direction from the Sun.
@@ 5 Simulated constraints
Three scenarios are presented in Figure 6. The first one (top left), assumes the array would be exclusively devoted to MBA occultations. In this case, the distance between telescopes is set at $D=2$ km. There is no change in $\sigma_{M}$ at larger $D$, since the majority of occulters are smaller than 2 km and the number of observed occultations becomes independent of $D$. The second scenario (top right) assumes an exclusive Jovian Trojan survey, and sets the distance between telescopes to $D=5$ km. Larger $D$ is favored for the Trojan survey, since they are twice as far as MBAs, making the smallest ones detectable by LSST larger than for MBAs. The third scenario (bottom) assumes a joint MBA/Trojan survey, with $D=5$ km again since the Trojans benefit and the MBA observations are not harmed by increased $D$.
@@
To assess the feasibility of performing observations at or close to the predicted event frequency, we assume that a single measurement uses $\sim$ 2 minutes of telescope time. The optimal configuration is therefore limited at $\sim$ 6000 events $\times$ telescopes per hour. Our simulations recorded hourly rates of successful observations (on nights with clear sky) varying in the 2000–3000 range. This means that, if exactly one telescope and time slot is required for each of 3000 successful events, the “subscription factor” would be $f\approx 0.5$ on the busiest nights, and if the events are randomly distributed over stations and time, a fraction $e^{-f}-(1-f)\approx 0.1$ would overlap another event and be lost. If, however, 2 or more time slots are needed for each success, either because the orbit uncertainty spans multiple telescopes, or the targets are small enough to slip between stations, then $f\approx 1$ or more, meaning $\approx 40\%$ or more of events would be lost to resource conflicts. This is most concerning at the initial stage of the survey, when target ephemerides are limited by LSST astrometry, and we expect a few failures before attaining the first successful observation of each target. Therefore, we would definitely benefit from array design options that push towards higher event frequency capabilities, and/or faster event cycles than 2 minutes. If the survey, before most objects get their first detection, is ultimately limited by the event load, we might need to extend survey duration to recover our predicted constraints.
@@ 6 Conclusions
The constraint on missing mass in the solar system would not be the only science benefit of the small-telescope linear array. Our survey would allow the study of other perturbations to asteroid orbits, constraining Yarkovsky acceleration parameters and detecting mutual deflections by close asteroid encounters. These data, therefore, would contribute to a better understanding of asteroid densities, albedos, and thermal properties. Such an array would also provide detailed shapes for $\sim 1800$ TNOs, characterizing the distribution of sizes, shapes, albedos, and contact-binary rates of TNOs in many different dynamical states. No other envisioned observation could yield such information for a large sample of pristine products of planetesimal formation.
@@ Appendix B Measurement errors on chord midpoints
where we have defined $B=n_{b}/n_{\star}$ as the background ratio. In the first equality, we have exploited the fact that the Fisher information in a Poisson distribution with a mean of $\lambda(q)$ for some parameter $q$ is equal to $(d\lambda/dq)^{2}/\lambda,$ and we can divide the light curve into statistically independent time segments of duration $\Delta t$ with expected mean count rates of $\lambda=n(t)\,\Delta t.$